This reaction shows the complete combustion of octane, C8H18, a component of gasoline.
2 C8H18(l) + 25 O2(g) 16 CO2(g) + 18 H2O(l)

(a) How many moles of O2 are needed to burn 1.00 mol of C8H18?

mol
(b) How many moles of CO2 are produced when 0.62 mol of C8H18 are burned?

mol
(c) How many grams of O2 are needed to burn 1.00 g of C8H18?
1
Expert's answer
2018-01-25T12:31:07-0500
(a):
Amount of O2: 25 / 2 = 12.5 mol
(b):
Amount of CO2: 0.62 ∙ (16 / 2) = 4.96 mol
(c):
Amount of C8H18: 1 / 114.23 = 0.008754 mol
Amount of O2: 0.008754 ∙ 12.5 = 0.109425 mol
Mass of O2: 0.109425 ∙ 32.00 = 3.5016 g
Answer:
(a): 12.5 mol
(b): 4.96 mol
(c): 3.5016 g